The two companies with top revenues in the Fortune 500 list for 2005 were Exxon Mobil and WlMart. Their revenues together totaled $656 billion. Wal-Mart revenues were $24 billion less than Exxon Mobil revenues. What were the revenues of each corporation?
Answer by tutorcecilia(2152)


Mobil + (Mobil-$24 billion) = $656 billion
x+(x-24)=656
2x-24=656
2x=656+24
2x=689
x=340
.
So,
Mobil = $340 billion
Wal-Mart = x-24=340-24=316
.
Checking:
340+316=656 [checks out]